# Fair Conformal Prediction for Incomplete Covariate Data

### v1

@@ 3.2 Efficient conformal prediction by localization
We now provide a new localized conformal prediction with missing covariates. Note that Assumption 2 is also crucial in the localized conformal prediction. We begin with finding the available cases for $M^{(n+1)}$ in the training and calibration set, i.e., ~$\tilde{\mathcal{I}}_{\text{train}}=\{i\in\mathcal{I}_{\text{train}}:M^{(i)}\preceq M^{(n+1)}\}$ and ~=$\tilde{\mathcal{I}}=\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\}$. Then, the scores can be written as
@@
5:Find available cases in training set and calibration set: ~$\tilde{\mathcal{I}}_{\text{train}}=\left\{i\in\mathcal{I}_{\text{train}}:M^{(i)}\preceq M^{(n+1)}\right\}$ and ~=$\tilde{\mathcal{I}}=\left\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\right\}$
@@ 4.1 Synthetic experiments
In synthetic experiments, data are generated following the Gaussian linear model described in Example 1. Specifically, the dimension of $X$ is set to 3, 5, and 8, respectively. The covariates are drawn from a normal distribution, $X\sim N(\mu,\Sigma)$, where $\mu=(1,\dots,1)^{T}$ and $\Sigma=\phi(1,\dots,1)^{T}(1,\dots,1)+(1-\phi)I_{d}$, with $\phi=0.8$. Gaussian noise is added, with $\varepsilon\sim N(0,1)$. The regression coefficient is set to $\beta=(1,2,-1,3,-0.5,-1,0.3,1.7)^{T}$. For dimensions $d=3$ and $d=5$, only the first $d$ covariates and corresponding parameters are used in the model.
@@
The training and calibration sets consist of 500 and 250 samples, respectively. For evaluating marginal coverage and interval length, the test set size is 2000. To assess the mask-conditional coverage, $\mathbb{P}\left\{Y\in\hat{C}_{\alpha}\left(X,m\right)\mid M=m\right\}$, and the interval length for each $m\in\{0,1\}^{d}$, we maintain a fixed number of observations per mask. This approach controls for variability in coverage and prediction interval length unrelated to $\mathbb{P}\{M=m\}$. As $d$ increases, the number of possible missing data patterns grows exponentially; for example, with $d=8$, there are $2552^{8}-1=$ 255 unique patterns. When $d=5$ or $d=8$ under MCAR and MNAR settings, we present coverage and interval length as a function of mask size, serving as a proxy for mask-conditional coverage. For each mask size, 100 observations are drawn according to the distribution of $M|\text{size}(M)$ in the test set. This evaluation is based on 50 replications, with the coverage and interval lengths averaged across experiments.
Tables 1-3 present the simulation results. Overall, all six methods achieve marginal validity. However, CQR-MDA-Nested appears conservative, with coverage levels significantly exceeding the nominal coverage level of $0.91-\alpha=$ 0.9. In terms of mask-conditional validity, CP and CQR do not meet this criterion. Specifically, CP?s prediction interval length remains invariant to the mask, and while CQR?s prediction interval length varies slightly with the mask, this variation is insufficient to achieve mask-conditional validity. The remaining methods, CQR-MDA-Exact, CQR-MDA-Nested, nexCP, and LCP, are mask-conditionally valid, with CQR-MDA-Nested again demonstrating conservativeness. Furthermore, the proposed nexCP and LCP are more efficient, yielding shorter prediction intervals than CQR-MDA-Exact and CQR-MDA-Nested. For instance, in Table 2, with the dimension of $X$ set to 5, both nexCP and LCP generally provide shorter intervals than CQR-MDA-Exact and CQR-MDA-Nested. Additionally, LCP shows slightly shorter intervals than nexCP, likely due to LCP?s ability to utilize quantitative information among covariates, whereas nexCP relies solely on qualitative information when constructing prediction intervals.
@@
CQR CQR-MDA-Exact CQR CQR-MDA-Exact CQR CQR-MDA-Exact
@@
CQR CQR-MDA-Exact CQR CQR-MDA-Exact CQR CQR-MDA-Exact
@@
CQR CQR-MDA-Exact CQR CQR-MDA-Exact CQR CQR-MDA-Exact
@@ 4.2 Semi-synthetic experiments
CQR CQR-MDA-Exact CQR CQR-MDA-Exact CQR CQR-MDA-Exact
@@ 5 Real data analysis
CQR CQR-MDA-Exact CQR CQR-MDA-Exact CQR CQR-MDA-Exact
@@ 6 Conclusion
The proof of the marginal validity is quite standard, the similar strategy can also be found in, e.g., [22, 4]. By i.i.d. observations and Assumption 1, the data points $(X^{(i)},M^{(i)},Y^{(i)}),i\in[n+1]$, are indeed exchangeable, and the regression and imputation algorithms treat these data points symmetrically. Thus, the absolute residuals $V^{(i)}=|Y^{(i)}-\hat{\mu}\circ\hat{\Phi}(X^{(i)},M^{(i)})|,i\in[n+1]$, are exchangeable.
@@
$\mathbb{P}\left\{Y^{(n+1)}\notin\hat{\mathcal{C}}_{n}\left(X^{(n+1)},M^{(n+1)}\right)\right\}$ =$=\mathbb{P}\{n+1\in S(V)\}=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbb{P}\{i\in S(V)\}$
@@
A.2 Example 1: Gaussian linear model.
The Gaussian pattern mixture model in (7) is
$(X_{1},X_{2})^{T}\mid M=m\sim N\left(\begin{pmatrix}\mu_{1}^{m}\\ \mu_{2}^{m}\end{pmatrix},\begin{pmatrix}\Sigma_{11}^{m}&\Sigma_{12}^{m}\\ \Sigma_{21}^{m}&\Sigma_{22}^{m}\end{pmatrix}\right),$
where $X_{1}$ is missing, and $X_{2}$ is observed. According to the conditional distribution of the multivariate distribution, we have
$X_{1}|(X_{2},M=m)\sim N(\mu_{1|2}^{m},\Sigma_{1|2}^{m}),$
where $\mu_{1|2}^{m}=\mu_{1}^{m}+\Sigma_{12}^{m}{\Sigma_{22}^{m}}^{-1}(X_{2}-\mu_{2}^{m}),\Sigma_{1|2}^{m}=\Sigma_{11}^{m}-\Sigma_{12}^{m}{\Sigma_{22}^{m}}^{-1}\Sigma_{21}^{m}$. Given that $Y=\beta^{T}X+\varepsilon$, with $\varepsilon\sim N(0,\sigma^{2})\perp\!\!\!\perp(X,M)$, the following holds:
[…]
By the fact that $(\beta_{1}^{T}X_{1})|(X_{2},M=m)\sim N(\beta_{1}^{T}\mu_{1|2}^{m},\beta_{1}^{T}\Sigma_{1|2}^{m}\beta_{1}),\varepsilon|(X_{2},M=m)\sim N(0,\sigma^{2})$, and $(\beta_{1}^{T}X_{1}\perp\!\!\!\perp\varepsilon)|(X_{2},M=m)$, we have
$Y|(X_{\text{obs}(m)},M=m):=Y|(X_{2},M=m)\sim N\left(\beta_{2}^{T}X_{2}+\beta_{1}^{T}\mu_{1\mid 2}^{m},\beta_{1}^{T}\Sigma_{1\mid 2}^{m}\beta_{1}+\sigma^{2}\right).$
A.3 Proof of Lemma 1.
@@
In the proposed nonexchangeable conformal prediction (Algorithm 1), we find the available cases for the mask of the test point, i.e., ~=$\tilde{\mathcal{I}}=\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\}$. As we focus on the mask-conditional validity, in the subsequent analysis, we will set $M^{(n+1)}=m$. Note that for $\tilde{m}\preceq m$, the distribution of $(X_{\text{obs}(\tilde{m})},\tilde{m},Y)$ can be written as
@@
By comparing equation (A.7) and (A.9), $P_{Y|(X_{\text{obs}(m)},M=\tilde{m})}=P_{Y|(X_{\text{obs}(m)},M=m)}$ by Assmption (2), and $P_{X_{\text{obs}(m)}|M=\tilde{m}}=P_{X_{\text{obs}(m)}|M=m}$ by the MCAR Assumption. Thus, there is no distribution shift between $(X_{\text{obs}(m)},m,Y)$ and $(X_{\text{obs}(m)},\tilde{m},Y)$, for $\tilde{m}\preceq m$. Therefore, $V^{(n+1)}=|Y^{(n+1)}-\hat{\mu}\circ\hat{\Phi}(X^{(n+1)},m)|$ and $\tilde{V}^{(i)}=|Y^{(i)}-\hat{\mu}\circ\hat{\Phi}(X^{(i)},m)|,i\in\tilde{\mathcal{I}}$ are exchangeable, so $\tilde{V}^{(i)}_{*},i\in[|\tilde{\mathcal{I}}|]$ and $V^{(n+1)}$ are exchangeable. There is no coverage loss by introducing fixed weights when the scores are indeed exchangeable [4]. Thus, for the fixed weights $\tilde{w}_{i},i\in[|\tilde{\mathcal{I}}|$ and $\tilde{w}_{n+1}$, the nonexchangeable conformal prediction in Algorithm 1 satisfyes
$\mathbb{P}\left\{Y^{(n+1)}\in\hat{\mathcal{C}}_{n}\left(X^{(n+1)},m\right)|M=m\right\}$ =$=\mathbb{P}\left\{V^{(n+1)}\leq Q_{1-\alpha}\left({\textstyle\sum_{i\in[|\tilde{\mathcal{I}}|]}}\tilde{w}_{i}\cdot\delta_{\tilde{V}_{*}^{(i)}}+\tilde{w}_{n+1}\cdot\delta_{+\infty}\right)\right\}$
@@
=$=\mathbb{P}\left\{V_{\alpha,h}\left(X^{(n+1)},m,Y^{(n+1)}\right)\leq Q_{1-\alpha}\left(\frac{1}{|\tilde{\mathcal{I}}|+1}\sum_{i\in\tilde{\mathcal{I}}}\delta_{V_{\alpha,h}^{(i)}}+\frac{1}{|\tilde{\mathcal{I}}|+1}\delta_{+\infty}\right)\right\},$
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 3.2 Efficient conformal prediction by localization
We now provide a new localized conformal prediction with missing covariates. Note that Assumption 2 is also crucial in the localized conformal prediction. We begin with finding the available cases for $M^{(n+1)}$ in the training and calibration set, i.e., $\tilde{\mathcal{I}}_{\text{train}}=\{i\in\mathcal{I}_{\text{train}}:M^{(i)}\preceq M^{(n+1)}\}$ and $\tilde{\mathcal{I}}=\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\}$. Then, the scores can be written as
@@
5:Find available cases in training set and calibration set: $\tilde{\mathcal{I}}_{\text{train}}=\left\{i\in\mathcal{I}_{\text{train}}:M^{(i)}\preceq M^{(n+1)}\right\}$ and $\tilde{\mathcal{I}}=\left\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\right\}$
@@ 4.1 Synthetic experiments
In synthetic experiments, We adopt a simulation design similar to that in [45], to enable direct comparison. Data are generated following the Gaussian linear model:
$Y=\beta^{T}X+\varepsilon.$ (11)
Specifically, the dimension of $X$ is set to 3, 5, and 8, respectively. The covariates are drawn from a normal distribution, $X\sim N(\mu,\Sigma)$, where $\mu=(1,\dots,1)^{T}$ and $\Sigma=\phi(1,\dots,1)^{T}(1,\dots,1)+(1-\phi)I_{d}$, with $\phi=0.8$. Gaussian noise is added, with $\varepsilon\sim N(0,1)$. The regression coefficient is set to $\beta=(1,2,-1,3,-0.5,-1,0.3,1.7)^{T}$. For dimensions $d=3$ and $d=5$, only the first $d$ covariates and corresponding parameters are used in the model.
@@
The training and calibration sets consist of 500 and 250 samples, respectively. For evaluating marginal coverage and interval length, the test set size is 2000. To assess the mask-conditional coverage, $\mathbb{P}\left\{Y\in\hat{C}_{\alpha}\left(X,m\right)\mid M=m\right\}$, and the interval length for each $m\in\{0,1\}^{d}$, we maintain a fixed number of observations per mask. This approach controls for variability in coverage and prediction interval length unrelated to $\mathbb{P}\{M=m\}$. As $d$ increases, the number of possible missing data patterns grows exponentially; for example, with $d=8$, there are $2^{8}-1=255$ unique patterns. When $d=5$ or $d=8$ under MCAR and MNAR settings, we present coverage and interval length as a function of mask size, serving as a proxy for mask-conditional coverage. For each mask size, 100 observations are drawn according to the distribution of $M|\text{size}(M)$ in the test set. This evaluation is based on 50 replications, with the coverage and interval lengths averaged across experiments.
Tables 1-3 present the simulation results. Overall, all six methods achieve marginal validity. However, CQR-MDA-Nested appears conservative, with coverage levels significantly exceeding the nominal coverage level of $1-\alpha=0.9$. In terms of mask-conditional validity, CP and CQR do not meet this criterion. Specifically, CP?s prediction interval length remains invariant to the mask, and while CQR?s prediction interval length varies slightly with the mask, this variation is insufficient to achieve mask-conditional validity. The remaining methods, CQR-MDA-Exact, CQR-MDA-Nested, nexCP, and LCP, are mask-conditionally valid, with CQR-MDA-Nested again demonstrating conservativeness. Furthermore, the proposed nexCP and LCP are more efficient, yielding shorter prediction intervals than CQR-MDA-Exact and CQR-MDA-Nested. For instance, in Table 2, with the dimension of $X$ set to 5, both nexCP and LCP generally provide shorter intervals than CQR-MDA-Exact and CQR-MDA-Nested. Additionally, LCP shows slightly shorter intervals than nexCP, likely due to LCP?s ability to utilize quantitative information among covariates, whereas nexCP relies solely on qualitative information when constructing prediction intervals.
@@
CQR CQR-MDA-Exact CQR-MDA-Nested CP nexCP LCP
@@
CQR CQR-MDA-Exact CQR-MDA-Nested CP nexCP LCP
@@
CQR CQR-MDA-Exact CQR-MDA-Nested CP nexCP LCP
@@ 4.2 Semi-synthetic experiments
CQR CQR-MDA-Exact CQR-MDA-Nested CP nexCP LCP
@@ 5 Real data analysis
CQR CQR-MDA-Exact CQR-MDA-Nested CP nexCP LCP
@@ 6 Conclusion
The proof of the marginal validity is quite standard, the similar strategy can also be found in, e.g., [26, 6, 45]. By i.i.d. observations and Assumption 1, the data points $(X^{(i)},M^{(i)},Y^{(i)}),i\in[n+1]$, are indeed exchangeable, and the regression and imputation algorithms treat these data points symmetrically. Thus, the absolute residuals $V^{(i)}=|Y^{(i)}-\hat{\mu}\circ\hat{\Phi}(X^{(i)},M^{(i)})|,i\in[n+1]$, are exchangeable.
@@
$\mathbb{P}\left\{Y^{(n+1)}\notin\hat{\mathcal{C}}_{n}\left(X^{(n+1)},M^{(n+1)}\right)\right\}$ $=\mathbb{P}\{n+1\in S(V)\}=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbb{P}\{i\in S(V)\}$
@@
A.2 Proof of Lemma 1.
@@
In the proposed nonexchangeable conformal prediction (Algorithm 1), we find the available cases for the mask of the test point, i.e., $\tilde{\mathcal{I}}=\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\}$. As we focus on the mask-conditional validity, in the subsequent analysis, we will set $M^{(n+1)}=m$. Note that for $\tilde{m}\preceq m$, the distribution of $(X_{\text{obs}(\tilde{m})},\tilde{m},Y)$ can be written as
@@
[…]
$\mathbb{P}\left\{Y^{(n+1)}\in\hat{\mathcal{C}}_{n}\left(X^{(n+1)},m\right)|M=m\right\}$ $=\mathbb{P}\left\{V^{(n+1)}\leq Q_{1-\alpha}\left({\textstyle\sum_{i\in[|\tilde{\mathcal{I}}|]}}\tilde{w}_{i}\cdot\delta_{\tilde{V}_{*}^{(i)}}+\tilde{w}_{n+1}\cdot\delta_{+\infty}\right)\right\}$
@@
$=\mathbb{P}\left\{V_{\alpha,h}\left(X^{(n+1)},m,Y^{(n+1)}\right)\leq Q_{1-\alpha}\left(\frac{1}{|\tilde{\mathcal{I}}|+1}\sum_{i\in\tilde{\mathcal{I}}}\delta_{V_{\alpha,h}^{(i)}}+\frac{1}{|\tilde{\mathcal{I}}|+1}\delta_{+\infty}\right)\right\},$
@@
By the definition of $\theta^{(i)}$, we have $s_{i}^{\theta^{(i)}}=0$. In addition, by the monotonicity of $s(v,F,\theta)$ with respect to $\theta$ and the definitional of $\theta^{*}$, that is, $\theta^{*}$ is less or equal to at least the $(1-\alpha)$-quantile of all $\theta^{(i)}$ and $\{+\infty\}$, we have at least $(1-\alpha)$-quantile of $s_{i}^{\theta^{*}}\leq 0$. So, one get $h(\theta^{*})\leq 0$. Therefore, for the fixed $\theta^{*}$ given ecent $E_{\tilde{v}}$, we have
@@
$\sup_{p\in\Sigma(\gamma,L)}\mathbb{E}\left[\int(\hat{p}_{h}(x)-p_{h}(x))^{2}dx\right]=O\left(\left(\frac{1}{n}\right)^{\frac{2\gamma}{2\gamma+k}}\right),$
@@
$\mathbb{E}\int\left[\hat{F}_{h}(V\mid X=x,M=m)-F_{h}(V|X=x,M=m)\right]^{2}dx$
$\leq\sup_{p\in\Sigma(\gamma,L)}\mathbb{E}\int(\hat{p}_{h}(x)-p(x))^{2}dx=O\left(\left(\frac{1}{n}\right)^{\frac{2\gamma}{2\gamma+k}}\right),$